OpenStudy (anonymous):

find f'(x): f(x) = x*sqrt(2-x^2)

OpenStudy (amistre64):

looks like a product and a chain wrapped together with some other basics thrown in for good measure

OpenStudy (amistre64):

f(x) = x*sqrt(2-x^2) f'(x) = (x*sqrt(2-x^2))' f'(x) = x'sqrt(2-x^2) + x*sqrt(2-x^2)' f'(x) = x'sqrt(2-x^2) + x*sqrt(2-x^2)' * (2-x^2)'

OpenStudy (savvy):

\[x \sqrt{2-x ^{2}} - (x ^{2}/\sqrt{2-x ^{2}})\]

OpenStudy (anonymous):

(2-2x^2)/sqrt(2-x^2) thanks i got it tho